# Rankers, Judges, and Assistants: Towards Understanding the Interplay of LLMs in Information Retrieval Evaluation

### v1

@@ Abstract
issue: 1issue: 2issue: 3issue: 4
1 Introduction
@@
Recent research has begun to explore some of these emerging issues. For example, studies have shown that LLMs can exhibit biases in their output, favoring LLM-generated content over human-generated ones [9], and perpetuating biases present in their training data [14, 20, 33]. Furthermore, LLM-based rating systems have been found to be susceptible to manipulation [2], may not accurately reflect human preferences [28], and demonstrate self-inconsistency [50]. Additionally, the phenomenon of “model collapse” has also been observed, where LLMs trained on synthetic data generated by other LLMs can lead to a degradation of quality and diversity in generated content [47].
@@
Within the IR research community, the use of LLMs for assessment is a subject of ongoing debate [13], with opinions ranging from complete rejection of LLMs for relevance assessment [48] to the assertion that they can fully replace human judgments [55]. Investigations have thus far focused on the agreement of LLM-generated ratings with human assessments [13, 55, 54] and the potential for LLMs to introduce biases in search results [9]. However, a comprehensive analysis of the implications of LLMs across the entire information ecosystem, from content creation with AI assistance to LLM-based reranking and LLM-based judges for evaluation, remains a critical gap in the current literature.
@@
This paper aims to advance our understanding of these issues by synthesizing prior research and, crucially, providing novel empirical evidence. We specifically focus on the novel challenge of understanding the effect LLM-based rankers and AI-powered content creation have on an LLM-based judge’s ability to accurately assess relevance. Prior work has separately noted the potential interaction between LLM-based rankers and judges [43, 13, 54, 30] (an interaction that has yet to be empirically investigated), while other initial work has explored the relationship between AI-powered content creation and rankers [9]. However, we argue that the complex interplay between each of these roles must be considered holistically to fully understand the potential implications of widespread adoption of LLM-based judges. We present initial results demonstrating the importance of this interconnected perspective, showcasing how the use of LLMs across the information lifecycle can influence the accuracy and potential biases of LLM judges.
@@
We start by considering the case of LLMs being used as both rankers and judges and present the first empirical demonstration of a significant bias of LLM judges towards LLM-based rankers. Novel to our approach is the examination of LLM judge performance via the use of oracle rankers, allowing for a controlled assessment of LLM judge behavior and discriminative ability. Using the TREC 2019 and 2020 Deep Learning track datasets, we conduct experiments that also compare different-sized LLM judges within the same model family. Our results reveal several key findings: (1) LLM judges are more lenient in their relevance assessments than human judges, confirming previous observations [55]; (2) LLM judges exhibit a significant bias towards LLM-based rankers, a phenomenon previously only hypothesized; and (3) LLM judges demonstrate limited ability to discern subtle, yet statistically significant, performance differences between systems. Additionally, we conduct a preliminary study into whether LLM judges demonstrate biases when they encounter AI-generated content. Contrary to some previously published findings [26, 37, 27], our experiments do not provide evidence of this bias, suggesting that deeper, more rigorous empirical evaluations are required to better understand this phenomenon.
@@
2 Background
Figure 1: LLM usage in modern information access systems.
@@ 2.1 LLMs as Rankers
In modern large-scale IR systems, a multi-stage retrieve-then-rerank pipeline has become a prominent approach, wherein an initial retrieval stage, often based on lexical matching or embedding-based methods, is followed by one or multiple reranking stages, utilizing more sophisticated models to refine the results. This reranking stage frequently employs LLMs, either fine-tuned for the task of ranking or utilized in a LLM prompting paradigm.
Work on fine-tuned rerankers focus on the supervised learning to rank problem [24, 40] by fine-tuning Pre-trained Language Models (PLMs) such as T5 [35, 64] or BERT [34]. The reranker can also be used in a pairwise fashion with two documents as the input [39]. More recently, prompting-based LLM rerankers have drawn much attention due to their strong off-the-shelf ranking performance, which can be roughly catgorized in to pointwise [19, 11, 45], pairwise [41], and listwise [51, 29] approaches, depending on the number of documents that scored or compared in the input prompt.
Dai et al. [9] present results suggesting an inherent bias in neural retrieval models toward LLM-generated texts. This source bias may stem from shared Transformer-based architectures and pretraining approaches, and can lead to “semantic shortcuts” during matching. Neural IR models are also shown to be vulnerable to adversarial attacks, such as keyword stuffing and content injection [38, 52].
2.2 LLMs as Judges
Early LLMs, such as BERT, have been utilized for measuring the distributional similarity of texts [60, 61] and for evaluating specific tasks via fine-tuning, including machine translation [63], text summarization [25], and question answering [31]. The arrival of generative LLMs, such as ChatGPT, have enabled various data labeling and annotation tasks [16]. The use of LLMs as surrogates for humans for evaluation, often referred to as “LLM-as-a-Judge” [62], now extends across virtually all natural language processing tasks, including text summarization and dialog response generation [17]. However, recent research increasingly demonstrates their limitations, such as favoring longer responses (length bias[12, 58] or content generated by similar models (self bias[27, 59].
@@ 2.2 LLMs as Judges
Our interest is specifically in the use of LLMs for relevance assessments in IR. MacAvaney and Soldaini [30] is among the first to employ LLMs for automatic relevance labeling. They specifically focus on a setting where a single known relevant document per query is available for evaluation and explore several one-shot approaches, and report high correlation with the ranking of systems from full TREC assessments. Faggioli et al. [13] present a spectrum of human-machine collaboration for producing relevance assessments, from AI assistance to fully automated judgments. They conduct a preliminary assessment of LLMs’ capabilities of relevance judgments on two TREC collections and report a fair agreement between human assessors and LLMs. They further report strong agreement in the relative ranking of systems using human vs. LLM relevance labels, but measures computed under the LLM judgments are found to be less sensitive than measures computed under human judgments. Thomas et al. [54] experiment with various prompt templates to improve quality and observe better agreement with the official labels on the TREC Robust 2004 collection than Faggioli et al. [13]. These improvements are attributed to both prompt design and the use of a more capable LLM (their in-house version of GPT-4). Thomas et al. [54] further share experiences on using LLMs for relevance assessment at Microsoft Bing, where LLMs have reportedly been used, in conjunction with expert human labelers, since late 2022. They utilize feedback from real searchers to select a prompt variant that agrees with this feedback (i.e., first-party ground truth) and find that LLMs can perform substantially better than third-party human judges. Importantly, the labels produced by the LLM are regularly checked by comparing them against labels by trained human assessors, ensuring quality assurance through manual verification. Upadhyay et al. [56] reproduce results from Thomas et al. [54], verifying their claims, and create an open-source implementation (UMBRELA). Most recently, LLMs are leveraged in the TREC 2024 Retrieval Augmented Generation (RAG) track for automatic relevance assessment [55]. Relative system rankings are found to correlate with those obtained using human judgments, even if human assessors apply stricter relevance criteria than LLMs [55]. The authors also experiment with various LLM-assisted labeling processes, such as using UMBRELA to pre-filter the pools or to suggest relevance labels that human judges can then post-edit, but find that those solutions “do not appear to have obvious tangible benefits over fully automatic processes” [55]. Clarke and Dietz [6] raise concerns about the claims made by Upadhyay et al. [55] and highlight how LLM-based judgments fail to demonstrate strong alignment with manual judgments for top-performing systems. They further present evidence that when evaluation is performed through a publicly known automatic process, such as UMBRELA, it can be subject to manipulation: a purpose-built submission to the TREC RAG track was placed as 5th using automatic evaluation while only 28th under manual assessment. Chen et al. [5] show that when performing relevance assessments in batches, the relevance levels of earlier documents in a batch influences the relevance judgments of subsequent documents, and that some LLMs are more affected by this so-called threshold priming effect than others. Alaofi et al. [2] compare various open-source and proprietary LLMs in labeling passages for relevance. Larger, i.e., more capable LLMs demonstrate comparable LLM-to-human agreement with human-to-human agreement (observed on TREC collections) on binary labels. When a graded scale is considered, only the largest commercial LLMs achieve the human-to-human agreement range. Larger LLMs also exhibit higher stability regardless of the prompt used. Interestingly, the authors demonstrate that most LLMs exhibit some degree of susceptibility to judging non-relevant documents as relevant if query words are inserted at random positions, simulating a keyword stuffing SEO strategy. Rahmani et al. [44] present a large-scale synthetic passage ranking collection, SnyDL, by extending the TREC 2019-2023 Deep Learning (DL) collections via LLM-generated labels. They observe a high agreement on system ordering between SynDL and the original TREC DL collections.
2.3 LLMs as Assistants
@@
There is a wide array of AI tools available to aid people with content creation. Focusing only on textual content, the spectrum ranges from basic grammar and spell checkers to advanced tools that generate full articles. Studies indicate that by late 2024, LLM assistance is detectable in a significant portion of various text domains, with estimates reaching up to 18% of financial consumer complaints and 24% in corporate press releases [21]. The use of powerful LLMs can lead to situations where it is unclear whether the content is primarily human-created with AI assistance or the other way around.
2.4 LLMs for Data Augmentation
@@
While not considered for this role in the current paper, LLMs are also used for data augmentation. For example, Dai et al. [10] use few-shot prompting to generate synthetic queries, while Bonifacio et al. [3] consider query generation in a full unsupervised setting. Soudani et al. [49] present a survey on synthetic dialogue data generation in open-domain, task-oriented, and information seeking dialogue systems.
3 Critical Issues with LLMs as Judges
@@ 3.1 Quality of Judgments
3.1.1 Validity and Discriminative Power
@@
  • Agreement on Individual Relevance Judgments: Several studies demonstrated that it is possible to use LLMs for relevance assessment and obtain performance comparable to TREC judges [13, 55] and notably better than crowd judges [54]. At the same time, it has also been observed that LLMs are more lenient when labeling a document relevant [55, 2], which leads to inflated evaluation scores. This leniency can lead to inflated evaluation scores, potentially masking subtle differences between systems.
@@
  • Agreement on System Rankings: A common approach to meta-evaluating LLM judges is to compare the relative ranking of retrieval systems based on LLM assessments with the ranking based on human-generated relevance judgments. This typically involves calculating the correlation between the two rankings, often using systems submitted to TREC tracks [30, 55, 13]. While high correlation is often interpreted as evidence of LLM judge validity, this approach has significant limitations.
@@
Discriminative Ability and the Limits of Correlation Even though several studies report on strong leaderboard correlation between human and LLM judgments, Clarke and Dietz [6] argue that Kendall’s $\tau$ is “less informative for assessing progress at the top of the leaderboard” and demonstrate that LLM-based assessments fail to reliably identify the best-performing systems. Further, Alaofi et al. [2] show that correlation-based meta-evaluation hides interesting failure patterns. A crucial, often overlooked, aspect is the disconnect between typical TREC evaluation setups and the needs of many practical IR scenarios. TREC evaluations often involve dozens of systems with widely varying approaches and performance levels. In contrast, practitioners often need to compare a small number of high-performing (state-of-the-art) systems or distinguish between subtle variations of a single system (e.g., in ablation studies). It remains an open question whether LLM judges possess the necessary sensitivity to reliably detect small but meaningful performance differences in such scenarios. Indeed, achieving high correlation is inherently easier with a larger and more diverse set of systems; simply including more systems with varying performance levels can artificially inflate correlation, even if the LLM judge struggles to differentiate between the top contenders. It thus remains an open question: Can LLM judges reliably distinguish between high-performing systems with small, but meaningful, performance differences?
3.1.2 Reliability and Reproducibility
@@
Beyond validity, a critical concern for LLM-based evaluation is the reliability and reproducibility of the judgments. Even if an LLM demonstrates a reasonable level of agreement with human judgments on average, its utility as a judge is undermined if its assessments are highly sensitive to seemingly minor variations in setup or input. Indeed, existing research demonstrates that LLM judgments can be significantly influenced by factors such as the choice of LLM [13, 2, 5], the specific wording and structure of the prompt [2, 54], and even the order in which documents are judged [5]. This variability raises concerns about the reliability of results obtained with a single LLM.
@@
The Impact of Model Choice A recurring theme in the literature is that more powerful LLMs (typically larger models with more parameters and trained on larger datasets) tend to exhibit better performance and consistency as judges [2]. This raises a crucial, but largely unexplored, question: To what extent would the conclusions of a study change if a more (or less) powerful LLM were used as the judge? This sensitivity to model choice has not been systematically investigated, particularly in the context of comparing high-performing systems where subtle differences matter.
3.2 Vulnerability to Bias and Manipulation

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730348isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Information retrieval
1. Introduction
@@
Recent research has begun to explore some of these emerging issues. For example, studies have shown that LLMs can exhibit biases in their output, favoring LLM-generated content over human-generated ones (Dai et al., 2024b), and perpetuating biases present in their training data (Gallegos et al., 2024; Liang et al., 2021; Navigli et al., 2023). Furthermore, LLM-based rating systems have been found to be susceptible to manipulation (Alaofi et al., 2024), may not accurately reflect human preferences (Liu et al., 2024c), and demonstrate self-inconsistency (Stureborg et al., 2024). Additionally, the phenomenon of “model collapse” has also been observed, where LLMs trained on synthetic data generated by other LLMs can lead to a degradation of quality and diversity in generated content (Shumailov et al., 2024).
@@
Within the IR research community, the use of LLMs for assessment is a subject of ongoing debate (Faggioli et al., 2023), with opinions ranging from complete rejection of LLMs for relevance assessment (Soboroff, 2025) to the assertion that they can fully replace human judgments (Upadhyay et al., 2024a). Investigations have thus far focused on the agreement of LLM-generated ratings with human assessments (Faggioli et al., 2023; Upadhyay et al., 2024a; Thomas et al., 2024) and the potential for LLMs to introduce biases in search results (Dai et al., 2024b). However, a comprehensive analysis of the implications of LLMs across the entire information ecosystem, from content creation with AI assistance to LLM-based reranking and LLM-based judges for evaluation, remains a critical gap in the current literature.
@@
This paper aims to advance our understanding of these issues by synthesizing prior research and, crucially, providing novel empirical evidence. We specifically focus on the novel challenge of understanding the effect LLM-based rankers and AI-powered content creation have on an LLM-based judge’s ability to accurately assess relevance. Prior work has separately noted the potential interaction between LLM-based rankers and judges (Rahmani et al., 2024; Faggioli et al., 2023; Thomas et al., 2024; MacAvaney and Soldaini, 2023) (an interaction that has yet to be empirically investigated), while other initial work has explored the relationship between AI-powered content creation and rankers (Dai et al., 2024b). However, we argue that the complex interplay between each of these roles must be considered holistically to fully understand the potential implications of widespread adoption of LLM-based judges. We present initial results demonstrating the importance of this interconnected perspective, showcasing how the use of LLMs across the information lifecycle can influence the accuracy and potential biases of LLM judges.
@@
We start by considering the case of LLMs being used as both rankers and judges and present the first empirical demonstration of a significant bias of LLM judges towards LLM-based rankers. Novel to our approach is the examination of LLM judge performance via the use of oracle rankers, allowing for a controlled assessment of LLM judge behavior and discriminative ability. Using the TREC 2019 and 2020 Deep Learning track datasets, we conduct experiments that also compare different-sized LLM judges within the same model family. Our results reveal several key findings: (1) LLM judges are more lenient in their relevance assessments than human judges, confirming previous observations (Upadhyay et al., 2024a); (2) LLM judges exhibit a significant bias towards LLM-based rankers, a phenomenon previously only hypothesized; and (3) LLM judges demonstrate limited ability to discern subtle, yet statistically significant, performance differences between systems. Additionally, we conduct a preliminary study into whether LLM judges demonstrate biases when they encounter AI-generated content. Contrary to some previously published findings (Liu et al., 2023; Panickssery et al., 2024; Liu et al., 2024b), our experiments do not provide evidence of this bias, suggesting that deeper, more rigorous empirical evaluations are required to better understand this phenomenon.
@@
2. Background
Figure 1. LLM usage in modern information access systems.
@@ 2. Background
In modern large-scale IR systems, a multi-stage retrieve-then-rerank pipeline has become a prominent approach, wherein an initial retrieval stage, often based on lexical matching or embedding-based methods, is followed by one or multiple reranking stages, utilizing more sophisticated models to refine the results. This reranking stage frequently employs LLMs, either fine-tuned for the task of ranking (Nogueira et al., 2020; Zhuang et al., 2023; Nogueira and Cho, 2019; Pradeep et al., 2021) or via prompting in a pointwise (Liang et al., 2023; Drozdov et al., 2023; Sachan et al., 2022), pairwise (Qin et al., 2024), or listwise (Sun et al., 2023; Ma et al., 2023) fashion. Dai et al. (2024b) present results suggesting an inherent bias in neural retrieval models toward LLM-generated texts. This source bias may stem from shared Transformer-based architectures and pretraining approaches, and can lead to “semantic shortcuts” during matching. Neural IR models are also shown to be vulnerable to adversarial attacks, such as keyword stuffing and content injection (Parry et al., 2024; Tamber and Lin, 2025).
LLMs as Judges
@@
Early LLMs, such as BERT, have been utilized for measuring the distributional similarity of texts (Zhang et al., 2020; Zhao et al., 2019) and for evaluating specific tasks via fine-tuning, including machine translation (Zhu et al., 2020), text summarization (Liu and Lapata, 2019), and question answering (McCarley et al., 2019). The arrival of generative LLMs, such as ChatGPT, have enabled various data labeling and annotation tasks (Gilardi et al., 2023). The use of LLMs as surrogates for humans for evaluation, often referred to as “LLM-as-a-Judge” (Zheng et al., 2023), now extends across virtually all natural language processing tasks, including text summarization and dialog response generation (Gu et al., 2024). However, recent research increasingly demonstrates their limitations, such as favoring longer responses (length bias(Dubois et al., 2024; Wang et al., 2023) or content generated by similar models (self bias(Liu et al., 2024b; Xu et al., 2024). Our interest is specifically in the use of LLMs for relevance assessments in IR. MacAvaney and Soldaini (2023) is among the first to employ LLMs for automatic relevance labeling. They specifically focus on a setting where a single known relevant document per query is available for evaluation and explore several one-shot approaches. Faggioli et al. (2023) present a spectrum of human-machine collaboration for producing relevance assessments, from AI assistance to fully automated judgments. They conduct a preliminary assessment of LLMs’ capabilities of relevance judgments on two TREC collections and report a fair agreement between human assessors and LLMs. Thomas et al. (2024) experiment with various prompt templates to improve quality and observe better agreement with the official TREC labels than Faggioli et al. (2023). These improvements are attributed to both prompt design and the use of a more capable LLM. Thomas et al. (2024) further share experiences on using LLMs for relevance assessment at Microsoft Bing, where LLMs have reportedly been used, in conjunction with expert human labelers, since late 2022. Upadhyay et al. (2024b) reproduce results from Thomas et al. (2024), verifying their claims, and create an open-source implementation (UMBRELA). Most recently, LLMs are leveraged in the TREC 2024 Retrieval Augmented Generation (RAG) track for automatic relevance assessment (Upadhyay et al., 2024a). Relative system rankings are found to correlate with those obtained using human judgments, even if human assessors apply stricter relevance criteria than LLMs (Upadhyay et al., 2024a). The authors also experiment with various LLM-assisted labeling processes, such as using UMBRELA to pre-filter the pools or to suggest relevance labels that human judges can then post-edit, but find that those solutions “do not appear to have obvious tangible benefits over fully automatic processes” (Upadhyay et al., 2024a). Clarke and Dietz (2024) raise concerns about the claims made by Upadhyay et al. (2024a) and highlight how LLM-based judgments fail to demonstrate strong alignment with manual judgments for top-performing systems. They further present evidence that when evaluation is performed through a publicly known automatic process, such as UMBRELA, it can be subject to manipulation. Chen et al. (2024a) show that when performing relevance assessments in batches, the relevance levels of earlier documents in a batch influences the relevance judgments of subsequent documents, and that some LLMs are more affected by this so-called threshold priming effect than others. Alaofi et al. (2024) compare various open-source and proprietary LLMs in labeling passages for relevance. They demonstrate that most LLMs exhibit some degree of susceptibility to judging non-relevant documents as relevant if query words are inserted at random positions, simulating a keyword stuffing SEO strategy. Rahmani et al. (2025) present a large-scale synthetic passage ranking collection, SnyDL, by extending the TREC 2019-2023 Deep Learning collections via LLM-generated labels, and observe a high agreement on system ordering.
LLMs as Assistants
@@
There is a wide array of AI tools available to aid people with content creation. Focusing only on textual content, the spectrum ranges from basic grammar and spell checkers to advanced tools that generate full articles. Studies indicate that by late 2024, LLM assistance is detectable in a significant portion of various text domains, with estimates reaching up to 18% of financial consumer complaints and 24% in corporate press releases (Liang et al., 2025). The use of powerful LLMs can lead to situations where it is unclear whether the content is primarily human-created with AI assistance or the other way around.
LLMs for Data Augmentation
@@
While not considered for this role in the current paper, LLMs are also used for data augmentation. For example, Dai et al. (2023) use few-shot prompting to generate synthetic queries, while Bonifacio et al. (2022) consider query generation in a full unsupervised setting. Soudani et al. (2024) present a survey on synthetic dialogue data generation in open-domain, task-oriented, and information seeking dialogue systems. The use of LLM-generated data brings forth new challenges in bias and unfairness, potentially affecting the reliability of IR systems (Dai et al., 2024a).
3. Critical Issues with LLMs as Judges
@@ 3.1. Quality of Judgments
3.1.1. Validity and Discriminative Power
@@
  • Agreement on Individual Relevance Judgments: Several studies demonstrated that it is possible to use LLMs for relevance assessment and obtain performance comparable to TREC judges (Faggioli et al., 2023; Upadhyay et al., 2024a) and notably better than crowd judges (Thomas et al., 2024). At the same time, it has also been observed that LLMs are more lenient when labeling a document relevant (Upadhyay et al., 2024a; Alaofi et al., 2024), which leads to inflated evaluation scores. This leniency can lead to inflated evaluation scores, potentially masking subtle differences between systems.
@@
  • Agreement on System Rankings: A common approach to meta-evaluating LLM judges is to compare the relative ranking of retrieval systems based on LLM assessments with the ranking based on human-generated relevance judgments. This typically involves calculating the correlation between the two rankings, often using systems submitted to TREC tracks (MacAvaney and Soldaini, 2023; Upadhyay et al., 2024a; Faggioli et al., 2023). While high correlation is often interpreted as evidence of LLM judge validity, this approach has significant limitations.
@@
Issue #1: Discriminative Ability and the Limits of Correlation Even though several studies report on strong leaderboard correlation between human and LLM judgments, Clarke and Dietz (2024) argue that Kendall’s $\tau$ is “less informative for assessing progress at the top of the leaderboard” and demonstrate that LLM-based assessments fail to reliably identify the best-performing systems. Further, Alaofi et al. (2024) show that correlation-based meta-evaluation hides interesting failure patterns. A crucial, often overlooked, aspect is the disconnect between typical TREC evaluation setups and the needs of many practical IR scenarios. TREC evaluations often involve dozens of systems with widely varying approaches and performance levels. In contrast, practitioners often need to compare a small number of high-performing (state-of-the-art) systems or distinguish between subtle variations of a single system (e.g., in ablation studies). It remains an open question whether LLM judges possess the necessary sensitivity to reliably detect small but meaningful performance differences in such scenarios. Indeed, achieving high correlation is inherently easier with a larger and more diverse set of systems; simply including more systems with varying performance levels can artificially inflate correlation, even if the LLM judge struggles to differentiate between the top contenders. It thus remains an open question: Can LLM judges reliably distinguish between high-performing systems with small, but meaningful, performance differences?
3.1.2. Reliability and Reproducibility
@@
Beyond validity, a critical concern for LLM-based evaluation is the reliability and reproducibility of the judgments. Even if an LLM demonstrates a reasonable level of agreement with human judgments on average, its utility as a judge is undermined if its assessments are highly sensitive to seemingly minor variations in setup or input. Indeed, existing research demonstrates that LLM judgments can be significantly influenced by factors such as the choice of LLM (Faggioli et al., 2023; Alaofi et al., 2024; Chen et al., 2024a), the specific wording and structure of the prompt (Alaofi et al., 2024; Thomas et al., 2024), and even the order in which documents are judged (Chen et al., 2024a). This variability raises concerns about the reliability of results obtained with a single LLM.
@@
Issue #2: The Impact of Model Choice A recurring theme in the literature is that more powerful LLMs (typically larger models with more parameters and trained on larger datasets) tend to exhibit better performance and consistency as judges (Alaofi et al., 2024). This raises a crucial, but largely unexplored, question: To what extent would the conclusions of a study change if a more (or less) powerful LLM were used as the judge? This sensitivity to model choice has not been systematically investigated, particularly in the context of comparing high-performing systems where subtle differences matter.
3.2. Vulnerability to Bias and Manipulation